OpenStudy (anonymous):

subtract.. -5√192-2√243-8√108

OpenStudy (dumbcow):

Simplify the radicals by factoring out perfect squares: \[\sqrt{192} = \sqrt{4}\sqrt{48}=\sqrt{4}\sqrt{16}\sqrt{3} = 2*4*\sqrt{3} = 8\sqrt{3}\] Now try the other ones